A 25 mL of 6M sodium hydroxide are diluted to 1.3L solution. What is the resulting concentration?
1
Expert's answer
2013-01-17T10:50:03-0500
Molarity is calculated due to the following formula:

Cm = n/V
n(NaOH) = Cm*V
n(NaOH) = 6 mol/L * 0.025 L = 0.15 mol

Total volume:

V(solution) = 1300 + 25 = 1325 ml = 1.325 L

Concentration:

Cm(NaOH) = n/V
Cm(NaOH) = 0.15 mol/1.325L = 0.1132 M

Answer: 0.1132 M